mathslover (mathslover):

Differentiation Problem.... (5) Find the Derivative of \(\tan^{-1} \left(\cfrac{\sqrt{1+x^2} -1}{x} \right) \) w.r.t \(\cos^{-1} \left(\sqrt{\cfrac{1 + \sqrt{1+x^2}}{2\sqrt{1+x^2}}}\right) \)

mathslover (mathslover):

In order to give a brief idea about this... This question is from the topic - Differentiatin of Parametric Functions.

mathslover (mathslover):

And I'm not sure from where should I start dealing with it.

mathslover (mathslover):

@Callisto

OpenStudy (anonymous):

will start like this \( (\tan^{-1} ( g(x) ))' = \tan'^{-1} g(x) . g(x)' \)

OpenStudy (anonymous):

mmm chain rule

mathslover (mathslover):

Okay... tbh, I was not in the mood to try that. But, I will try this now with the help of ur suggestion... 1 minute

OpenStudy (anonymous):

ok shall i continue with u ? or u wanna do it by urself ?

mathslover (mathslover):

I will prefer that you continue with me... as I don't have proper plan in the mind to do it myself.

OpenStudy (anonymous):

ill give u the mood power :O Goooooooooooooooo 

OpenStudy (anonymous):

ok one thing i have to ask w.r.t cos.... means u wanna the answer with cos -1 ?

mathslover (mathslover):

lol...

mathslover (mathslover):

It is more like f(x) w.r.t g(x)

OpenStudy (anonymous):

hhh idk if that gonna work burt leet give it a try :)

OpenStudy (anonymous):

\(\large (\tan^{-1} ( g(x) ))' = \tan'^{-1} g(x) . g(x)' \) =\(\large = \frac {1}{1+g(x)} . g'(x) \)

OpenStudy (anonymous):

or u wanna proof that tan^-1 g(x)= 1/1+x ??

OpenStudy (anonymous):

idk why Os is laging for me -,,-

OpenStudy (vishweshshrimali5):

Put, x = \(\tan {\theta}\)

OpenStudy (vishweshshrimali5):

Then, \[\tan^{-1} (\cfrac{\sqrt{1+x^2}-1}{x}) = \tan^{-1} (\cfrac{\sec{\theta} - 1}{\tan{\theta}})\]

OpenStudy (vishweshshrimali5):

\[= \tan^{-1}(\cfrac{1-\cos{\theta}}{\sin{\theta}})\]

OpenStudy (vishweshshrimali5):

Put, cos \(\theta\) = 1-2\(sin^2{\theta}\)

mathslover (mathslover):

= theta/2 , right?

OpenStudy (vishweshshrimali5):

*theta/2 in sine

OpenStudy (vishweshshrimali5):

Yes

mathslover (mathslover):

Right, so, the next function is also theta/2 ?

mathslover (mathslover):

And thus : d(f(x))/d(g(x)) = 1 ? Right?

mathslover (mathslover):

f(x) = tan^{-1} ... g(x) = cos^{-1} ...

OpenStudy (vishweshshrimali5):

Yes

mathslover (mathslover):

Got it.. Thanks a lot @vishweshshrimali5 . And thanks @BSwan .

OpenStudy (anonymous):

:D

mathslover (mathslover):

;)

OpenStudy (anonymous):

*head hurts* yup....time to go play a game to entertain my small attention span..